OpenStudy (anonymous):

two objects attract each other with gravitational force of 25 newtons from a given distances. if the distances between the two objects is reduced by a factor of 5, what is the changed force of attraction between them?

OpenStudy (anonymous):

F'/F=(Gm1m2/r'^2)/(Gm1m2/r^2) F'/F=r^2/r'^2 F'=(r/r')^2*F F'=(5)^2*25

OpenStudy (anonymous):

force between them will increse by the factor of 25

OpenStudy (anonymous):

125?